OpenStudy (anonymous):

Write a function rule that represents each sentence. 13. t is 4 more than the product of 7 and s. 14. The ratio of a to 5 equals b. 15. y is half of x plus 10. 16. d is 8 less than the quotient of b and 4.

OpenStudy (anonymous):

13. t= 7s+4

OpenStudy (anonymous):

15. y= (x/2)+10

OpenStudy (anonymous):

13.t=7s+4 14.a/5=b 15.y=x/2+10

OpenStudy (anonymous):

16. d= (b/4)-8